A two-digit number is 1 more than eight times the sum of the digits. The ones digit is 3 less than the tens digit. What is the number?
Answer by dkppathak(439)

A two-digit number is 1 more than eight times the sum of the digits. The ones digit is 3 less than the tens digit. What is the number?
let unit digit is X and tens digit is Y
number =10 multiplied by tens digit +unit place by using this formula
number will be =10 Y + x as per given conditions
10Y +X = 8(x +Y)+1
10Y+X=8x+8y+1
10y+X-8x-8y=1
2Y -7x=1 (1)
by second conditions
X =Y-3
Y-X = 3 (2) by multiplying 2
2Y-2X=6 (2)
2Y-7X =1(1) by subtraction
5x =5
X=1
by substitution X=1 in equation 2
Y-x =3
y-1 =3
Y=3+1 =4
tens place is 4 and unit place is 1
number will be 41